# Tighter Confidence Bounds for Sequential Kernel Regression

### v1

@@ 4.2 Confidence Sequences
We now describe how the tail bound in (8) can be used to construct confidence sequences for $f^{*}$. Let $\boldsymbol{f}_{t}=[f(x_{1}),\dots,f(x_{t})]^{\top}$ denote the values of an arbitrary function $f\in\mathcal{H}$ at the points $x_{1},\dots,x_{t}$. Since, with probability at least $1-\delta$, (8) holds for all $t\geq 1$, the sets of functions $f\in\mathcal{H}$ that (at each $t$) satisfy both (8) and $\left\|f\right\|_{\mathcal{H}}\leq B$ form a $1-\delta$ confidence sequence for $f^{*}$.
Lemma 4.2 (Confidence sequence).
For any $c>0$ and $\delta\in(0,1]$, it holds with probability at least $1-\delta$ that for all $t\geq 1$ simultaneously, the function $f^{*}$ lies in the set
$\mathcal{F}_{t}=\bigg{\{}f\in\mathcal{H}\,\bigg{|}\,$ $\|\boldsymbol{f}_{t}-\boldsymbol{y}_{t}\|_{2}\leq R_{t},~{}\left\|f\right\|_{\mathcal{H}}\leq B\bigg{\}}.$
The two constraints that define $\mathcal{F}_{t}$ are both ellipsoids in $\mathcal{H}$, meaning that $\mathcal{F}_{t}$ is an intersection of two ellipsoids. In contrast, most existing confidence sets for functions in RKHS’s (e.g. (Abbasi-Yadkori, 2012; Chowdhury & Gopalan, 2017)) are single ellipsoids centred at the kernel ridge estimate $\mu_{\alpha,t}$ (see Eq. (2)), for some regularisation parameter $\alpha>0$. It turns out that by taking a weighted sum of the constraints in $\mathcal{F}_{t}$, we can obtain a single-ellipsoid confidence set ~$\widetilde{\mathcal{F}}_{t}\supseteq\mathcal{F}_{t}$, which is also centred at the kernel ridge estimate.
@@
For any $c>0$, any $\alpha>0$ and any $\delta\in(0,1]$, it holds with probability at least $1-\delta$ that for all $t\geq 1$ simultaneously, $f^{*}$ lies in the set
~$\widetilde{\mathcal{F}}_{t}=\left\{f\in\mathcal{H}\,\bigg{|}\,\left\|\left(\boldsymbol{V}_{t}+\alpha\boldsymbol{I}_{\mathcal{H}}\right)^{1/2}(f-\mu_{\alpha,t})\right\|_{\mathcal{H}}\leq\widetilde{R}_{\alpha,t}\right\},$
@@
The proof (in App. B.2) involves rearranging the inequality
@@
into the constraint that defines ~$\widetilde{\mathcal{F}}_{t}$. This larger confidence sequence can be more easily compared to those in (Abbasi-Yadkori, 2012; Chowdhury & Gopalan, 2017), cf. Sec. 6.1.
@@ 4.3 Implicit Confidence Bounds
The optimisation variable in (9) is a function in the possibly infinite-dimensional RKHS $\mathcal{H}$, which makes it difficult to directly compute the numerical solution of (9). We now describe two finite-dimensional reformulations of (9).
As (9) contains a constraint on the RKHS norm of $f$, the (generalised) representer theorem (Kimeldorf & Wahba, 1971; Schölkopf et al., 2001) can be applied. This means that the solution of (9) can be expressed as a finite linear combination of the form $\boldsymbol{k}_{t+1}(x)^{\top}\boldsymbol{w}$, where $\boldsymbol{k}_{t+1}(x):=[k(x,x_{1}),\dots,k(x,x_{t}),k(x,x)]^{\top}$. When we substitute the functional form $f(\cdot)=\boldsymbol{k}_{t+1}(\cdot)^{\top}\boldsymbol{w}$ into the objective and both constraints, we obtain the finite-dimensional convex program in Thm. 4.4. (see App. B.3).
Theorem 4.4 (Representer theorem for UCB computation).
@@
Eq. (10) is a $(t+1)$-dimensional second order cone program, which can be solved efficiently using conic solvers from, for example, the CVXPY library (Diamond & Boyd, 2016). To illustrate the simplicity of solving (10) with such libraries, we provide sample code in App. E.1. For our second reformulation, we focus on the Lagrangian dual problem associated with (9).
Theorem 4.5 (Dual problem for UCB computation).
The dual problem associated with (9) is
@@
See (2) and (3) for the definitions of $\mu_{\eta_{2}/\eta_{1},t}(x)$ and $\rho_{\eta_{2}/\eta_{1},t}^{2}(x)$. We provide a proof in App. B.4 and focus here on the consequences of this result. By weak duality, the solution of (11) is always an upper bound on the solution of the primal problem in (9). Whenever there exists an $f$ that lies in the interior of the ellipsoids that define $\mathcal{F}_{t}$, as is usually the case, then strong duality holds and the solutions of (9) and (11) are actually equal. In App. B.4, we show that (11) can be further reduced to a one-dimensional optimisation problem. Using the substitution $\eta_{2}=\alpha\eta_{1}$ (for $\alpha>0$), the solution of (11) is equal to
$\min_{\alpha>0}\left\{\mu_{\alpha,t}(x)+\frac{\widetilde{R}_{\alpha,t}}{\sqrt{\alpha}}\rho_{\alpha,t}(x)\right\}.$ (12)
The objective function in (12) is not always convex in $\alpha$, though empirical evidence suggests that it is quasiconvex in $\alpha$, and could therefore be solved using the bisection method. Since $\mu_{\alpha,t}(x)$ is the predictive mean of a standard Gaussian process (GP) posterior and $\rho_{\alpha,t}^{2}(x)$ is its predictive variance, the objective function in (12) closely resembles GP-based confidence bounds (e.g. (Srinivas et al., 2010)). We can therefore interpret our confidence bound $\mathrm{UCB}_{\mathcal{F}_{t}}(x)$ as a GP-based confidence bound, which is minimised with respect to the regularisation parameter.
@@ 4.4 Explicit Confidence Bounds
In this section, we focus on explicit confidence bounds. These confidence bounds are mainly useful for deriving worst-case regret bounds with explicit dependence on $T$ (see Section 6.2), but are also cheaper to compute than the implicit confidence bounds in the previous subsection. From (12) in the previous subsection, we can immediately see that for any fixed $\alpha>0$, we have
$\mathrm{UCB}_{\mathcal{F}_{t}}(x)~{}\leq~{}\mu_{\alpha,t}(x)+\frac{\widetilde{R}_{\alpha,t}}{\sqrt{\alpha}}\rho_{\alpha,t}(x).$ (13)
It turns out that this upper bound on $\mathrm{UCB}_{\mathcal{F}_{t}}(x)$ is equal to the confidence bound $\mathrm{UCB}_{\widetilde{\mathcal{F}}_{t}}(x)$, where ~$\widetilde{\mathcal{F}}_{t}$ is the confidence set from Lemma 4.3.
Corollary 4.6 (Analytic UCB).
@@
$\mathrm{UCB}_{\widetilde{\mathcal{F}}_{t}}(x)~{}\leq~{}\mu_{\alpha,t}(x)+\frac{\widetilde{R}_{\alpha,t}}{\sqrt{\alpha}}\rho_{\alpha,t}(x).$ (14)
This statement follows from the proof of Thm. 11. In particular, one can show that the RHS of (14) is the solution of the dual problem associated with $\max_{f\in\widetilde{\mathcal{F}}_{t}}\{f(x)\}$. Again, when ~$\widetilde{\mathcal{F}}_{t}$ has an interior point (as is usually the case), the inequality in (14) is actually an equality. The inequality in Cor. 4.6 also follows from Cauchy-Schwarz (App. B.5).
@@ 4.5 Confidence Bound Maximisation
We now discuss how our upper confidence bounds can be maximised w.r.t. $x\in\mathcal{X}$, which is an important consideration for kernel bandits (see Sec. 5). If $\mathcal{X}$ has finite (and not too large) cardinality, then $\mathrm{UCB}_{\mathcal{F}_{t}}(x)$ (and upper bounds on it) can easily be maximised exactly.
If $\mathcal{X}$ is a continuous subset of $\mathbb{R}^{d}$, then maximising $\mathrm{UCB}_{\mathcal{F}_{t}}(x)$ or any of our upper bounds over $x$ exactly is intractable in general, though other similar confidence bounds have the same limitation. There are at least two options for approximate maximisation. One option is to use gradient-based methods as in (Flynn et al., 2023). Gradients (w.r.t. $x$) of the implicit confidence bounds in (10), (11) and (12) can be computed numerically using differentiable convex optimsation methods (Agrawal et al., 2019) or automatic implicit differentiation methods (Look et al., 2020; Blondel et al., 2022). Gradient-based maximisation works well in practice, even though there are no rigorous approximation guarantees as $\mathrm{UCB}_{\mathcal{F}_{t}}(x)$ is (in general) a non-concave function of $x$. The second option is to discretise $\mathcal{X}$ and optimise over the finite discretisation (see e.g. Sec. 3.1 of (Li & Scarlett, 2022)). When the kernel function $k(x,x^{\prime})$ is Lipschitz (as the RBF and Matérn kernels are), then the discretisation error can be controlled. However, the grid size for a sufficiently fine discretisation grows exponentially in the dimension $d$ of $\mathcal{X}$, so this quickly becomes impractical.
5 Application To Kernel Bandits
@@ 5.1 Stochastic Kernel Bandits
A learner plays a game over a sequence of $T$ rounds, where $T$ may not be known in advance. In each round $t$, the learner observes an action set $\mathcal{X}_{t}\subseteq\mathcal{X}$ and must choose an action $x_{t}\in\mathcal{X}_{t}$. The learner then receives a reward $y_{t}=f^{*}(x_{t})+\epsilon_{t}$. The reward function $f^{*}:\mathcal{X}\to\mathbb{R}$ is an unknown function in an RKHS. As before, we assume that $f^{*}$ has bounded RKHS norm, i.e. $\left\|f^{*}\right\|_{\mathcal{H}}\leq B$, and that $\epsilon_{1},\epsilon_{2},\dots,\epsilon_{T}$ are conditionally $\sigma$-sub-Gaussian.
The goal of the learner is to choose a sequence of actions that minimises cumulative regret, which is the difference between the total expected reward of the learner and the optimal strategy. For a single round, we define the regret as $r_{t}=f^{*}(x_{t}^{*})-f^{*}(x_{t})$, where $x_{t}^{*}=\operatornamewithlimits{argmax}_{x\in\mathcal{X}_{t}}\{f^{*}(x)\}$. After $T$ rounds, the cumulative regret is $r_{1:T}=\sum_{t=1}^{T}r_{t}$.
5.2 Kernel UCB Meta Algorithm
In Algorithm 1, we describe a general recipe for a kernel bandit algorithm, which is based on a sequence of upper confidence bounds $u_{0},u_{1},u_{2},\dots$ for the reward function $f^{*}$. In each round $t$, we use all previous observations $\{(x_{s},y_{s})\}_{s=1}^{t-1}$ to successively construct the UCB $u_{t-1}$, and then select the action that maximises $u_{t-1}$.
@@ 5.2 Kernel UCB Meta Algorithm
For each $t$, we must be able to compute $u_{t}$ using the data $x_{1},y_{1},\dots,x_{t},y_{t}$. Kernel bandit algorithms of this form have appeared under different names, such as GP-UCB (Srinivas et al., 2010) and KernelUCB (Valko et al., 2013), based on the confidence bounds that they use. We therefore use the names KernelUCB and GP-UCB synonymously. When we use the confidence bounds in Thm. 3.11 of (Abbasi-Yadkori, 2012), we call Algorithm 1 AY-GP-UCB. If we use the confidence bounds in Thm. 2 of (Chowdhury & Gopalan, 2017), then Algorithm 1 is the Improved GP-UCB algorithm (IGP-UCB).
@@ 5.3 Kernel CMM-UCB
For our first variant of KernelUCB, we use our exact confidence bound $\mathrm{UCB}_{\mathcal{F}_{t}}(x)$. In particular, we set $u_{t}(x)$ to be the numerical solution of (10). The resulting algorithm, which we call Kernel Convex Martingale Mixture UCB (Kernel CMM-UCB), is a kernelised version of the CMM-UCB algorithm for linear bandits (Flynn et al., 2023).
@@ 5.4 Kernel DMM-UCB
$u_{t}(x)=\min_{\alpha\in A}\left\{\mu_{\alpha,t}(x)+\frac{\widetilde{R}_{\alpha,t}}{\sqrt{\alpha}}\rho_{\alpha,t}(x)\right\}.$
We call the resulting bandit algorithm Kernel Dual Martingale Mixture UCB (Kernel DMM-UCB). Because we only optimise $\alpha$ over the grid $A$, this upper confidence bound is looser than the exact confidence bound used by Kernel CMM-UCB. In return, the cost of computing this UCB is only a factor of $|A|$ more than the cost of computing the analytic UCB in Cor. 4.6.

### v2

@@ 4.2 Confidence Sequences
We now describe how the inequality in (8) can be used to construct confidence sequences for $f^{\star}$. Let $\boldsymbol{f}_{t}=[f(x_{1}),\dots,f(x_{t})]^{\top}$ denote the values of an arbitrary function $f\in\mathcal{H}$ at the points $x_{1},\dots,x_{t}$. Since, with probability at least $1-\delta$, (8) holds for all $t\geq 1$, the sets of functions that (at each $t$) satisfy both (8) and $\left\|f\right\|_{\mathcal{H}}\leq B$ form a $1-\delta$ confidence sequence for $f^{\star}$.
Lemma 4.2.
For any $c>0$ and $\delta\in(0,1]$, it holds with probability at least $1-\delta$ that for all $t\geq 1$ simultaneously, the function $f^{\star}$ lies in the set
$\mathcal{F}_{t}=\{f\in\mathcal{H}\,:\,$ $\|\boldsymbol{f}_{t}-\boldsymbol{y}_{t}\|_{2}\leq R_{t},~{}\left\|f\right\|_{\mathcal{H}}\leq B\}.$
The two constraints that define $\mathcal{F}_{t}$ are both ellipsoids in $\mathcal{H}$, meaning that $\mathcal{F}_{t}$ is the intersection of two ellipsoids. In contrast, most existing confidence sets for functions in RKHS’s (e.g. (Abbasi-Yadkori,, 2012; Chowdhury and Gopalan,, 2017)) are single ellipsoids centred at the kernel ridge estimate $\mu_{\alpha,t}$ (see Eq. (2)), for some regularisation parameter $\alpha>0$. It turns out that by taking a weighted sum of the constraints in $\mathcal{F}_{t}$, we can obtain a single-ellipsoid confidence set ~$\widetilde{\mathcal{F}}_{t}\supseteq\mathcal{F}_{t}$, which is also centred at the kernel ridge estimate.
@@
For any $c>0$, any $\alpha>0$ and any $\delta\in(0,1]$, it holds with probability at least $1-\delta$ that for all $t\geq 1$ simultaneously, $f^{\star}$ lies in the set
~$\widetilde{\mathcal{F}}_{t}=\{f\in\mathcal{H}\,:\,\|(\boldsymbol{V}_{t}+\alpha\boldsymbol{I}_{\mathcal{H}})^{1/2}(f-\mu_{\alpha,t})\|_{\mathcal{H}}\leq\widetilde{R}_{\alpha,t}\},$
@@
The proof (in App. B.2) involves rearranging
@@
into the constraint that defines ~$\widetilde{\mathcal{F}}_{t}$. This confidence sequence can be more easily compared to those in (Abbasi-Yadkori,, 2012; Chowdhury and Gopalan,, 2017).
@@ 4.3 Implicit Confidence Bounds
The optimisation variable in (9) is a function in the possibly infinite-dimensional RKHS $\mathcal{H}$. However, as (9) contains a constraint on the RKHS norm of $f$, the (generalised) representer theorem (Kimeldorf and Wahba,, 1971; Schölkopf et al.,, 2001) can be applied. This means that the optimising function $f$ in (9) can be expressed as a finite linear combination of the form $\boldsymbol{k}_{t+1}(x)^{\top}\boldsymbol{w}$, where $\boldsymbol{k}_{t+1}(x):=[k(x,x_{1}),\dots,k(x,x_{t}),k(x,x)]^{\top}$. When we substitute the functional form $f(\cdot)=\boldsymbol{k}_{t+1}(\cdot)^{\top}\boldsymbol{w}$ into the objective and both constraints, we obtain the finite-dimensional convex program in Thm. 4.4. (see App. B.3).
Theorem 4.4.
@@
Eq. (10) is a $(t+1)$-dimensional second-order cone program. Using interior point methods, the solution can be computed to $\epsilon$ accuracy in $\mathcal{O}(t^{3}\mathrm{polylog}(1/\epsilon))$ steps (see Sec. 10.2 of Nemirovski,, 2004). To obtain cheaper confidence bounds, we consider the Lagrangian dual problem associated with (9). We derive the dual in App. B.4 and give a short sketch here. The primal in (9) can be written as $\max_{f\in\mathcal{H}}\min_{\eta_{1},\eta_{2}>0}L(f,\eta_{1},\eta_{2})$, where
$L(f,\eta_{1},\eta_{2})=f(x)+\eta_{1}(R_{t}^{2}-\|\boldsymbol{f}_{t}-\boldsymbol{y}_{t}\|_{2}^{2})+\eta_{2}(B^{2}-\|f\|_{\mathcal{H}}^{2}),$
is the Lagrangian and $\eta_{1},\eta_{2}$ are the Lagrange multipliers. The dual problem $\min_{\eta_{1},\eta_{2}>0}\max_{f\in\mathcal{H}}L(f,\eta_{1},\eta_{2})$ is obtained by swapping the $\min$ and the $\max$. Since the Lagrangian is concave and quadratic in $f$, the inner maximisation in the dual can be solved analytically, which gives the following expression for the dual.
Theorem 4.5.
The dual of (9) is
@@
See (2) and (3) for the definitions of $\mu_{\eta_{2}/\eta_{1},t}(x)$ and $\rho_{\eta_{2}/\eta_{1},t}^{2}(x)$. By weak duality, the solution of (11) is always an upper bound on the solution of the primal problem in (9). Whenever there exists an $f$ that lies in the interior of the ellipsoids that define $\mathcal{F}_{t}$, as is usually the case, then strong duality holds and the solutions of (9) and (11) are actually equal. In App. B.4, we show that (11) can be further reduced to a one-dimensional optimisation problem. Using the substitution $\eta_{2}=\alpha\eta_{1}$ (for $\alpha>0$), the solution of (11) is equal to
$\min_{\alpha>0}\left\{\mu_{\alpha,t}(x)+\tfrac{1}{\sqrt{\alpha}}\widetilde{R}_{\alpha,t}\rho_{\alpha,t}(x)\right\}.$ (12)
The objective function in (12) is not always convex in $\alpha$, though empirical evidence suggests that it is quasiconvex in $\alpha$, and could therefore be solved using the bisection method.
@@ 4.4 Explicit Confidence Bounds
In this section, we focus on explicit confidence bounds. These are mainly useful for deriving worst-case regret bounds with explicit dependence on $T$ (see Sec. 6.2). From Eq. (12), for any fixed $\alpha>0$, we have
$\mathrm{UCB}_{\mathcal{F}_{t}}(x)~{}\leq~{}\mu_{\alpha,t}(x)+\tfrac{1}{\sqrt{\alpha}}\widetilde{R}_{\alpha,t}\rho_{\alpha,t}(x).$ (13)
It turns out that this upper bound on $\mathrm{UCB}_{\mathcal{F}_{t}}(x)$ is equal to $\mathrm{UCB}_{\widetilde{\mathcal{F}}_{t}}(x)$, where ~$\widetilde{\mathcal{F}}_{t}$ is the set from Lem. 4.3.
Corollary 4.6.
@@
$\mathrm{UCB}_{\widetilde{\mathcal{F}}_{t}}(x)~{}\leq~{}\mu_{\alpha,t}(x)+\tfrac{1}{\sqrt{\alpha}}\widetilde{R}_{\alpha,t}\rho_{\alpha,t}(x).$ (14)
This follows from the proof of Thm. 11. In particular, one can show that the RHS of (14) is the solution of the dual associated with $\max_{f\in\widetilde{\mathcal{F}}_{t}}\{f(x)\}$. Again, when ~$\widetilde{\mathcal{F}}_{t}$ has an interior point (as is usually the case), the inequality in (14) is actually an equality.
@@ 4.5 Confidence Bound Maximisation
We now discuss how our upper confidence bounds can be maximised w.r.t. $x\in\mathcal{X}$, which is an important consideration for kernel bandits (see Sec. 5). If $\mathcal{X}$ has finite (and not too large) cardinality, then $\mathrm{UCB}_{\mathcal{F}_{t}}(x)$ (and upper bounds on it) can be maximised exactly.
If $\mathcal{X}$ is a continuous subset of $\mathbb{R}^{d}$, maximising $\mathrm{UCB}_{\mathcal{F}_{t}}(x)$ over $\mathcal{X}$ exactly is intractable in general, though other similar confidence bounds have the same limitation. There are at least two options for approximate maximisation. One option is to use gradient-based methods, as in (Flynn et al.,, 2023). Gradients (w.r.t. $x$) of the confidence bounds in (10), (11) and (12) can be computed using differentiable convex optimisation methods (Agrawal et al.,, 2019) or automatic implicit differentiation methods (Look et al.,, 2020; Blondel et al.,, 2022). Gradient-based maximisation works well in practice, though there are no rigorous approximation guarantees as $\mathrm{UCB}_{\mathcal{F}_{t}}(x)$ is (in general) a non-concave function of $x$. A second option is to discretise $\mathcal{X}$ and optimise over the discretisation (see e.g. Sec. 3.1 of (Li and Scarlett,, 2022)). If the kernel function $k(x,x^{\prime})$ is Lipschitz (as the RBF and Matérn kernels are), then the discretisation error can be controlled. However, the size of a sufficiently fine discretisation grows exponentially in the dimension $d$, so this quickly becomes impractical.
5 KERNEL BANDITS
@@ 5.1 Stochastic Kernel Bandits
A learner plays a game over a sequence of $T$ rounds. In each round $t$, the learner observes an action set $\mathcal{X}_{t}\subseteq\mathcal{X}$ and must choose an action $x_{t}\in\mathcal{X}_{t}$. The learner then receives a reward $y_{t}=f^{\star}(x_{t})+\epsilon_{t}$. The reward function $f^{\star}:\mathcal{X}\to\mathbb{R}$ is an unknown function in an RKHS. As before, we assume that $f^{\star}$ has bounded norm, i.e. $\left\|f^{\star}\right\|_{\mathcal{H}}\leq B$, and that (in the noisy setting) $\epsilon_{1},\epsilon_{2},\dots,\epsilon_{T}$ are conditionally $\sigma$-sub-Gaussian.
The goal of the learner is to choose a sequence of actions that minimises the cumulative regret, which is the difference between the total expected reward of the learner and the optimal strategy. The regret in round $t$ is $r_{t}=f^{\star}(x_{t}^{\star})-f^{\star}(x_{t})$, where $x_{t}^{\star}=\operatornamewithlimits{argmax}_{x\in\mathcal{X}_{t}}\{f^{\star}(x)\}$. After $T$ rounds, the cumulative regret is $r_{1:T}=\sum_{t=1}^{T}r_{t}$.
5.2 Kernel Bandit Algorithms
In Algorithm 1, we present a recipe for a kernel bandit algorithm that uses confidence bounds to select actions.
@@ 5.2 Kernel Bandit Algorithms
In each round $t$, we use all previous observations $\{(x_{s},y_{s})\}_{s=1}^{t-1}$ to construct the UCB $u_{t-1}$, and then select the action that maximises $u_{t-1}$. Kernel bandit algorithms of this form have appeared under different names, such as GP-UCB (Srinivas et al.,, 2010) and KernelUCB (Valko et al.,, 2013), based on the confidence bounds that they use. We therefore use the names KernelUCB and GP-UCB synonymously. When we use the confidence bounds in Thm. 3.11 of (Abbasi-Yadkori,, 2012), we call Algorithm 1 AY-GP-UCB. If we use the confidence bounds in Thm. 2 of (Chowdhury and Gopalan,, 2017), then Algorithm 1 is the Improved GP-UCB algorithm (IGP-UCB).
@@ 5.3 Kernel CMM-UCB
For our first variant of KernelUCB, we use our exact confidence bound $\mathrm{UCB}_{\mathcal{F}_{t}}(x)$. In particular, we set $u_{t}(x)$ to be the numerical solution of (10). The resulting algorithm is called Kernel Convex Martingale Mixture UCB (Kernel CMM-UCB).
@@ 5.4 Kernel DMM-UCB
$u_{t}(x)=\min_{\alpha\in A}\left\{\mu_{\alpha,t}(x)+\tfrac{1}{\sqrt{\alpha}}\widetilde{R}_{\alpha,t}\rho_{\alpha,t}(x)\right\}.$
We call the resulting algorithm Kernel Dual Martingale Mixture UCB (Kernel DMM-UCB). Because we only optimise $\alpha$ over the grid $A$, this upper confidence bound is looser than the exact confidence bound used by Kernel CMM-UCB. In return, the cost of computing this UCB is only a factor of $|A|$ more than the cost of computing the analytic UCB in Cor. 4.6.